A uniform ladder 10m long and weighing 59N rets against a smooth vertical wall. If the ladder is just on the verge of slipping when it makes a 50 degree angle with the ground, find the coefficient of static friction between the ladder and the ground.
1
Expert's answer
2020-04-01T10:20:49-0400

As per the given question,

Length of the uniform ladder(L)=10m

Weight of the ladder =59N

Angle between the ladder and the ground ="50^\circ"

Let the reaction on the bottom of the ladder "N_1" and reaction on the wall "N_2"

Now, taking the torque about the center of the ladder

"N_1\times\dfrac{L}{2} \sin 50-\mu N_1\dfrac{L}{2}\cos 50=N_2\times \dfrac{L}{2}\cos 50"

"N_2=N_1\tan 50-\mu N_1--------(i)"


Now, balancing horizontal force

"N_2=\mu N_1-------(ii)"

From equation (i) and (ii)

"2\mu N_1=N_1\tan 50"

"\mu=\dfrac{\tan 50}{2}=0.59"

"\mu =0.59"